Can someone tell me if I have done this problem right?
(x+5)(x-4)=0
(x^2+5)(x+4)=0
x+5=0 or x+4=0
x=5 or x=4

Answer by rapaljer(4671)

If this is your problem:
(x+5)(x-4)=0
Then:
x+5=0 or x-4=0
x=-5 or x=4

R^2 at SCC